OpenStudy (anonymous):

Please help! Find the parametric representation of the curve x^2 + y^2 = 25 and z=(1/pi)arctan(x/y) needs to be in the form: x = x(t); y = y(t), z = z(t).

hero (hero):

\[z = \frac{1}{\pi} \times \arctan(\frac{x}{y})\]

hero (hero):

I'm wondering if that's what you mean

OpenStudy (anonymous):

yes it is :)

hero (hero):

I'll be back soon hopefully

OpenStudy (anonymous):

thanks!

hero (hero):

Okay, I am done, just need to post this

hero (hero):

If you have any questions about my solution, feel free to ask.

OpenStudy (anonymous):

thanks a lot!

hero (hero):

yw :D

hero (hero):

Sorry Mya, I can't be like you and use Latex.

OpenStudy (anonymous):

so if we use sin^2 t +cos^2 t =1 couldn't it also be y=5sint and x=cost and wouldn't this be a differrnt answer?